State whether each of the following statements is true or false. The unit of the rate constant for a first order reaction is M-2 * time-1 Tries 0/3 The unit of the rate constant for a third order reaction is M-1 * time-1 Tries 0/3 The reaction order that corresponds to the unit M-1/2 * s-1 is 1/2 Tries 0/3 The reaction order that corresponds to the unit year-1 is -1
Answer #1

Note:- Unit for n​​​​​​th order reaction will be given as follows.

Unit of k = M(1-n).s​​​​​​-1

or

Unit of k = M(1-n).(time)-1

where, time = second or minutes or hour or day or year etc

n = order of reaction

Part(1)

Unit of first order reaction

n = 1

So, Unit of k for first order reaction = M​​​​​​1 -1 .s​​​​​​-1

Unit of k for first order reaction = M​​​​​​0​​​.s-1

unit of k for first order reaction = s​​​​​​-1

So,

Answer = false

Part(2) unit of k for third order reaction

n =3

So,

Unit of k for third order reaction = M(1-3).s​​​​​​-1

Unit of k for third order reaction = M​​​​​​-2​.s-1

Answer = False

Part(3)

Unit of k = M​​​​​​-1/2.s-1

Comparing with equation number (1)

1-n = -1/2

n = 1+1/2

n = 3/2

(Hence , order of reaction = 3/2)

Answer = False

part(4)

unit of k = year​​​-1

Or

Unit of k = M​​​​​​0​​.year-1

Comparing with

Unit of k = M​​​​​​1-n.​​year-1

1-n = 0

n = 1

( Hence, Order or reaction = 1)

Answer = False